### v1

# Effect of Domain Generalization Techniques in Low-Resource Systems

Abstract
Machine learning models typically assume that training and test data follow the same distribution, an assumption that often fails in real-world scenarios due to distribution shifts. This issue is especially pronounced in low-resource settings, where data scarcity and limited domain diversity hinder robust generalization. Domain generalization (DG) approaches address this challenge by learning features that remain invariant across domains, often using causal mechanisms to improve model robustness. In this study, we examine two distinct causal DG techniques in low-resource natural language tasks. First, we investigate a causal data augmentation (CDA) approach that automatically generates counterfactual examples to improve robustness to spurious correlations. We apply this method to sentiment classification on the NaijaSenti Twitter corpus, expanding the training data with semantically equivalent paraphrases to simulate controlled distribution shifts. Second, we explore an invariant causal representation learning (ICRL) approach using the DINER framework, originally proposed for debiasing aspect-based sentiment analysis. We adapt DINER to a multilingual setting by constructing Afri-SemEval, a new benchmark derived from the SemEval-2014 Task 4 dataset via translation, covering 17 African languages. Through comprehensive experiments, we compare the effectiveness of these techniques in mitigating domain shifts and improving model performance. Our findings demonstrate that both approaches enhance robustness to unseen domains: counterfactual data augmentation yields consistent cross-domain accuracy gains in sentiment classification, while causal representation learning with DINER improves out-of-distribution performance in multilingual sentiment analysis, albeit with varying gains across languages. These results highlight the complementary strengths and trade-offs of data-level vs. representation-level causal interventions in low-resource AI settings.
Keywords Causal representation $\cdot$ Low-resource $\cdot$ Data augmentation $\cdot$ Language translation $\cdot$ LLMs
1 Introduction
Machine learning (ML) models have achieved remarkable success in diverse natural language processing (NLP) applications, yet their effectiveness largely depends on the assumption that training and test data are independently and identically distributed (i.i.d.). In practice, data distributions often shift across time, geography, and linguistic context, leading to significant performance degradation on unseen domains. This problem is especially acute in low-resource settings, where data scarcity and linguistic variation limit the ability of models to generalize robustly[1]
Domain generalization (DG) seeks to mitigate this challenge by learning representations that remain invariant across multiple source domains without direct access to the target distribution. Within DG, causal inference provides a principled foundation: by focusing on stable cause–and–effect relationships, models can identify invariant mechanisms that hold under interventions or domain shifts[2]. Causal modeling aims to capture the underlying generative processes common to all domains, thereby improving robustness when faced with distributional changes.
In this study, we examine two complementary causal DG approaches for low-resource sentiment analysis in African languages:
  • Invariance via Causal Data Augmentation (CDA): We leverage GPT-4o-mini to generate semantically equivalent paraphrases that preserve sentiment polarity while introducing lexical and syntactic diversity. This process expands the training data and simulates a controlled domain shift (from original domain $D_{1}$ to augmented domain $D_{2}$) under the constraint $P(y\mid x^{\prime})=P(y\mid x)$. Using the Yoruba and Igbo subsets of the NaijaSenti dataset[3], we assess how exposure to paraphrastic variation improves generalization and reduces the invariance gap $\Delta_{\text{inv}}$.
  • Invariance via Causal Representation Learning (CRL): We adopt the DINER framework[4], which models sentiment prediction as a Structural Causal Model (SCM) with explicit backdoor adjustments and counterfactual reasoning to isolate causal sentiment paths. To extend DINER to low-resource African contexts, we introduce Afri-SemEval, a multilingual benchmark translated from the SemEval-2014 Task 4 ABSA dataset[5], covering 17 African languages. Each translated instance retains the original review, target aspect, and sentiment polarity, allowing us to evaluate model invariance across a diverse set of languages and linguistic conditions.
By integrating these two frameworks, CDA for data-level causal perturbations and CRL for representation-level causal learning, we evaluate whether causal interventions at both levels can jointly enhance robustness under distributional shifts. Empirical results across Yoruba, Igbo, and the 17 Afri-SemEval languages show consistent gains in cross-domain accuracy and reduced sensitivity to spurious correlations. In particular, augmenting training data with counterfactual paraphrases yields more invariant performance between original and shifted distributions, while causal representation learning improves out-of-distribution generalization in multilingual sentiment analysis, with the magnitude of improvement varying based on linguistic and data characteristics.
Our contributions are threefold:
  • We propose a unified experimental design that combines data-level and representation-level causal interventions for low-resource African sentiment analysis.
  • We introduce Afri-SemEval, a multilingual extension of the SemEval-2014 ABSA dataset (translated into 17 African languages) to facilitate research on domain generalization in multilingual contexts.
  • We empirically demonstrate that incorporating causal mechanisms, through both data augmentation and causal representation learning, enhances generalization and robustness under distribution shifts in multilingual, low-resource NLP tasks.
2 Related Works
2.1 Domain Generalization and the Low-Resource Paradox
Recent comprehensive surveys reveal fundamental challenges in achieving OOD robustness for NLP systems [6, 7]. While domain adaptation techniques assume access to unlabeled target data, domain generalization confronts a more stringent constraint: models must extrapolate to completely unseen distributions using only source domain knowledge. Critically, the DomainBed evaluation demonstrated that carefully tuned empirical risk minimization often matches or exceeds specialized DG algorithms [8], suggesting that algorithmic complexity may not always translate to practical gains—a finding particularly relevant when computational resources are constrained.
For African languages, this theoretical challenge becomes acute. The NaijaSenti corpus [3] and AfriSenti benchmark [9] represent pioneering efforts, yet even these datasets contain only thousands of examples per language compared to millions available for high-resource languages. Community-driven documentation reveals that 88% of African languages lack basic computational resources [10], creating what we term a “low-resource paradox”: the very languages requiring the most robust generalization techniques have the least data to learn invariant representations from. Analysis of counterfactual data reveals that spurious correlations in small datasets can achieve training accuracies comparable to true causal features [11], explaining why models trained on limited data catastrophically fail when deployment conditions deviate even slightly from training distributions.
2.2 From Correlation to Causation: Theoretical Foundations
The causal revolution in machine learning offers a principled solution: if models learn causal mechanisms rather than statistical associations, they should maintain performance under distribution shifts [12, 13]. Comprehensive taxonomies of causal inference applications in NLP systematically categorize text as treatment, outcome, or confounder [14], yet most work focuses on high-resource English datasets, leaving open questions about scalability to low-resource, morphologically complex languages.
Two paradigms have emerged for implementing causal invariance. Invariant Risk Minimization penalizes representations where optimal classifiers vary across training environments [15], though subsequent analyses revealed optimization challenges and limited guarantees beyond linear settings [8]. More recently, explicit intervention methods construct task-specific Structural Causal Models to surgically remove spurious pathways [7, 2]. Our work bridges these paradigms by evaluating both data-level counterfactual generation and model-level causal debiasing in a unified low-resource experimental framework.
2.3 Counterfactual Augmentation and LLM-Based Data Generation
Counterfactual data augmentation, where annotators make minimal edits to flip labels while preserving coherence, demonstrates that models trained on counterfactually-augmented data exhibit 40-60% reduction in sensitivity to spurious features [11]. However, human annotation at scale remains prohibitive for low-resource languages. Active learning approaches with counterfactual generative models have achieved 18-20% robustness improvements using only 10% human-annotated counterfactuals [16]—a crucial finding for resource-constrained scenarios.
The emergence of instruction-tuned LLMs as data augmentation engines has been both promising and problematic. Recent experiments demonstrate that modern LLMs can expand 500 labeled examples to 5,000 synthetic samples with comparable downstream performance [17], but critically note that augmentation quality degrades substantially for low-resource languages like Tamil. This observation motivates our systematic evaluation of GPT-4o-mini [18] on Yoruba and Igbo—languages where paraphrase quality cannot be assumed but must be empirically validated. Unlike generic data augmentation techniques operating on token-level perturbations without semantic constraints [19], our approach explicitly enforces sentiment preservation through prompt engineering to maintain the counterfactual validity condition $P(y|x^{\prime})=P(y|x)$.
2.4 Causal Debiasing for Aspect-Based Sentiment Analysis
Aspect-Based Sentiment Analysis exposes fundamental weaknesses in neural architectures. The ARTS adversarial benchmark revealed that state-of-the-art ABSA models suffer up to 69.73% accuracy drops when aspect-sentiment correlations are manipulated [20], confirming that models exploit spurious statistical patterns [21, 22, 23] rather than genuine compositional understanding. While various debiasing strategies have been proposed [24, 25], most address single-variable inputs and fail to account for ABSA’s inherent multi-input structure.
The DINER framework represents a methodological advance by constructing an SCM distinguishing review-level confounding from aspect-level spurious correlation, applying backdoor adjustment to the former and counterfactual TIE estimation to the latter[4]. This hybrid strategy outperforms single-intervention baselines on English ABSA benchmarks. However, its applicability beyond English remains unexplored—a critical gap given that the SemEval-2014 Task 4 benchmark [5] exists only in English and a handful of European languages.
Our contribution addresses this gap through Afri-SemEval, translating the canonical ABSA benchmark into 17 typologically diverse African languages. This enables the first systematic evaluation of causal debiasing methods in multilingual low-resource contexts. Leveraging African-focused multilingual models with 17-language adaptive fine-tuning [26] and data-efficient pretraining [27], we investigate whether causal interventions transfer across linguistic typologies. The mixed results we observe illuminate a crucial limitation: translation-induced artifacts (word order changes, morphological mismatches, aspect-review misalignment) introduce unintended distribution shifts that interact with the intended cross-lingual domain shift [7]. This dual-shift phenomenon suggests that future causal DG methods must be robust not only to natural linguistic variation but also to the systematic noise inherent in cross-lingual transfer pipelines—a challenge distinct from the monolingual OOD scenarios dominating current literature.
3 Methodology
This section presents the methodological design adopted in this study, comprising two complementary experimental frameworks: (1)Invariance via Causal Data Augmentation and (2) Invariance via Causal Representation Learning.
3.1 Invariance via Causal Data Augmentation
This study investigates causal data augmentation as a strategy for improving sentiment classification in low-resource African languages through enhanced representation invariance. The approach is grounded in the principles of causal representation learning: if the sentiment label $y$ causally depends on the semantic content of a tweet $x$, then augmentations that perturb only non-causal linguistic features (e.g., surface form, style, or syntax) should preserve the conditional distribution $P(y|x)$. By systematically generating linguistically diverse yet semantically equivalent paraphrases, the model learns to focus on the underlying causal semantics rather than superficial correlations, leading to improved generalization across domains.
3.1.1 Data and Resources
We utilize the Yoruba and Igbo subsets of the NaijaSenti corpus[3], a multilingual sentiment-labeled Twitter dataset covering four Nigerian languages: Yoruba, Igbo, Hausa, and Nigerian Pidgin. Each instance is a tweet–label pair $(x_{i},y_{i})$, where $x_{i}$ represents the tweet and $y_{i}\in\{\text{positive},\text{negative},\text{neutral}\}$ is its sentiment label.
3.1.2 Training Configuration
Let $\mathcal{D}_{1}=\{(x_{i},y_{i})\}_{i=1}^{N}$ denote the original data distribution drawn from $P(X,Y)$. We define an augmentation function $f:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ that generates paraphrased examples $x^{\prime}=f(x)$ such that the sentiment semantics are preserved, i.e.,
$P(y|x^{\prime})=P(y|x).$ (1)
To achieve this, we employ the GPT-4o-mini model[18] with a carefully designed prompt that instructs the model to generate paraphrases that preserve sentiment polarity while introducing lexical and syntactic diversity. The prompt template used is:
"You have been provided with {language} tweet. Paraphrase ’{tweet}’ in a way that preserves its sentiment and meaning but changes its wording. Return only the paraphrased tweet."
This produces a new dataset $\mathcal{D}_{2}=\{(x^{\prime}_{i},y_{i})\}_{i=1}^{N}$ that shares semantic equivalence with $\mathcal{D}_{1}$ but differs in linguistic realization, thereby simulating a controlled domain shift:
$\begin{split}P_{\mathcal{D}_{2}}(X^{\prime})&\neq P_{\mathcal{D}_{1}}(X),\quad\text{while}\\ P_{\mathcal{D}_{2}}(Y|X^{\prime})&\approx P_{\mathcal{D}_{1}}(Y|X).\end{split}$ (2)
We train three distinct logistic regression classifiers for each language subset; (1)Original-only–trained on $\mathcal{D}_{1}$. (2) Paraphrased-only–trained on $\mathcal{D}_{2}$. (3) Combined–trained on $\mathcal{D}_{\text{comb}}=\mathcal{D}_{1}\cup\mathcal{D}_{2}$ (50-50 mix).
Each model is evaluated under a cross-domain testing scheme to measure generalization under distributional shift; Models trained on $\mathcal{D}_{1}$ are tested on $\mathcal{D}_{1}$, $\mathcal{D}_{2}$ and $\mathcal{D}_{\text{comb}}$, Models trained on $\mathcal{D}_{2}$ are tested on $\mathcal{D}_{2}$, $\mathcal{D}_{1}$ and $\mathcal{D}_{\text{comb}}$, Models trained on $\mathcal{D}_{\text{comb}}$ are tested on all three domains.
For each configuration, the model $h_{\theta}$ minimizes the empirical cross-entropy loss:
$\hat{\mathcal{L}}(\theta)=\frac{1}{|\mathcal{D}_{\text{train}}|}\sum_{(x,y)\in\mathcal{D}_{\text{train}}}\ell(h_{\theta}(g_{\psi}(\phi(x))),y),$ (3)
where $\phi$ represents text preprocessing (tokenization, normalization, stopword removal), and $g_{\psi}$ denotes the embedding function implemented using the paraphrase-multilingual-MiniLM-L12-v2[28] sentence transformer.
0ur Model performance was evaluated using accuracy computed over multiple domains, and invariance was quantified by measuring performance consistency across test domains:
$\Delta_{\text{inv}}=|\text{Acc}_{\mathcal{D}_{1}}-\text{Acc}_{\mathcal{D}_{2}}|,$ (4)
where smaller $\Delta_{\text{inv}}$ indicates stronger invariance to paraphrastic shifts. This design was adopted to enable a principled test of the causal invariance hypothesis, which posits that models trained with causal data augmentation should minimize $\Delta_{\text{inv}}$ and achieve superior generalization in low-resource settings.
3.2 Invariance via Causal Representation Learning
Invariance via Causal Representation Learning (CRL) during inference is one of the causal domain generalization techniques we aim to explore in low-resource settings, as outlined by Liu et al. [1]. Our goal is to examine how domain generalization (DG) can be effectively applied under such conditions. To achieve this, we adopt the DINER: Debiasing Aspect-based Sentiment Analysis with Multi-variable Causal Inference[4], a debiasing approach for the Aspect-Based Sentiment Analysis (ABSA) task. We further discuss DINER and our methodology in subsequent sections.
Figure 1 demonstrates how high-capacity neural networks can easily learn spurious correlations arising from annotation bias. These biases cause models to capture misleading dependencies rather than true causal relationships. Due to their high representational power, neural models are particularly vulnerable to such biases, often learning shortcuts instead of underlying latent features [21, 22, 23]. For instance, reversing the polarity of target aspects can result in up to a 20% drop in accuracy, highlighting the model’s reliance on these correlations. The figure also supports the observation that models frequently depend on spurious correlations between target and non-target aspects in sentiment datasets such as Pontiki et al. [5]. In the ’Original’ setting, both review-only and aspect-only models perform well; however, when these correlations are disrupted in the ’RevTgt’, ’RevNon’ and ’AddDiff’ settings, performance decreases or becomes balanced. This confirms that neural models often exploit dataset biases rather than learning genuine target-specific sentiment representations.
3.2.1 Debiasing in Aspect Review using DINER
To mitigate these effects, recent studies have proposed several debiasing strategies, including argumentation-based methods[19], reweighting-based training[24], and causal inference-based approaches[25, 1]. Among these, causal inference has attracted significant interest for its solid theoretical foundation and minimal disruption to existing learning paradigms.
The DINER framework[4] models ABSA as a Structural Causal Model (SCM) with five key variables: confounder $C$ (prior contextual knowledge), review $R$, aspect $A$, fused knowledge $K$ (derived from $R$ and $A$), and sentiment label $L$. The causal relations are defined as $R\rightarrow K\leftarrow A$, $K\rightarrow L$, $R\rightarrow L\leftarrow A$, and $C\rightarrow R,C\rightarrow L$. The causal structure is illustrated in Figure 2, showing the directed acyclic graph (DAG) that represents relationships among these variables. The objective of DINER is to isolate the causal path $K\rightarrow L$ while eliminating spurious correlations via $R\rightarrow L\leftarrow A$ and mitigating confounding through $C\rightarrow R$ and $C\rightarrow L$. For details on causal graph construction, please refer to Wu et al. [4].
Figure 1: Demonstration of spurious correlations learned from annotation bias ([4]).
Figure 2: Causal structure of the DINER framework ([4]).
To mitigate the indirect confounding between sentiment words in the review and the context, DINER employs a backdoor adjustment intervention that addresses the $C\rightarrow R\rightarrow L$ pathway. The intervention uses the do-calculus formulation:
$\begin{split}P(L|do(R))&=\sum_{c}P(L|R,C)P(C)\\ &=\sum_{c}\frac{P(L,R|C)P(C)}{P(R|C)}\end{split}$ (5)
The debiased review representation is computed as:
[…]
To address the direct correlation between aspect and label ($A\rightarrow L$), DINER employs counterfactual reasoning using the Total Indirect Effect (TIE):
$\begin{split}TIE_{a,r}&=TE_{a,r}-NDE_{r}-NDE_{a}+IE_{a,r}\\ &=TE_{a,r}-(NDE_{r}+NDE_{a})\quad(\text{since }IE_{a,r}=0)\end{split}$ (7)
Following Niu et al. [25], we define:
$\begin{split}TE_{a,r}&=L_{a,r,k}-L_{a^{*},r^{*},k^{*}}\\ NDE_{a}&=L_{a,r^{*},k^{*}}-L_{a^{*},r^{*},k^{*}}\\ NDE_{r}&=L_{a^{*},r,k^{*}}-L_{a^{*},r^{*},k^{*}}\end{split}$ (8)
Thus:
$\begin{split}TIE_{a,r}=L_{a,r,k}-L_{a^{*},r,k^{*}}\\ -L_{a,r^{*},k^{*}}+L_{a^{*},r^{*},k^{*}}\end{split}$ (9)
The final prediction is implemented through a model ensemble:
$\begin{split}L_{a,r,k}&=L(A\!=\!a,R\!=\!r^{\prime},K\!=\!k)\\ &=\Psi(\zeta_{a},\zeta_{r^{\prime}},\zeta_{k})\\ &=\zeta_{k}\!+\!\tanh(\zeta_{a})\!+\!\tanh(\zeta_{r^{\prime}})\end{split}$ (10)
Where $\zeta_{r^{\prime}}$ is the output of the debiased review-only branch, $\zeta_{a}$ is the output of the aspect-only branch, and $\zeta_{k}$ is the output of the fused knowledge branch.
3.2.2 Data
We used Afri-SemEval, a multilingual dataset created by translating the SemEval 2014 [5]) dataset into 17 African languages. The dataset spans diverse languages from multiple African language families, as shown in Table 1, [29].
Each translated instance preserves the original structure of SemEval, including the review text, target aspect term, and sentiment polarity (posistive, negative, neutral). Translation challenges included maintaining aspect-term alignment, preserving sentiment polarity, and capturing cultural nuances in sentiment expression.
Table 1: List of 17 African supported languages
S/N Language Code S/N Language Code
1 Afrikaans af 10 Kinyarwanda rw
2 Amharic am 11 Shona sn
3 Hausa ha 12 Somali so
4 Igbo ig 13 S. Sotho st
5 Lingala ln 14 Tswana tn
6 Luganda lug 15 Xhosa xh
7 N. Sotho nso 16 Yoruba yo
8 Chichewa nya 17 Zulu zu
9 Kirundi run
Each translated instance maintains the original SemEval-2014 format, comprising review text, target aspect, and sentiment polarity (positive, negative, or neutral). The translation process involved ensuring aspect alignment, polarity preservation, and accounting for linguistic and cultural nuances in sentiment expression.
3.3 African Language Adaptation
A major challenge in adapting DINER to African languages lies in aspect term alignment. The original framework assumes a direct mapping between aspect terms and their positions in the review, but translation often disrupts this due to word order variation (e.g., English SVO vs. Hausa SXV), morphological complexity (e.g., noun classes in Bantu languages), and differences in expression length (multiword to single-word mappings, and vice versa).
We address this with an alignment heuristic based on semantic similarity. Embeddings are computed for translated aspect terms and all n-grams in the review. The n-gram with the highest similarity to the aspect term is then selected as its aligned position for DINER processing. While effective for single-token aspects, this heuristic remains challenging for multi-token and non-contiguous aspect terms, which are common in many African languages.
4 Experiment and Result
4.1 Invariance via Causal Data Augmentation Results
All experiments were conducted on the Yoruba and Igbo subsets of the NaijaSenti dataset[3], following the causal data augmentation setup described in Section 3.1. For each language, three logistic regression classifiers were trained separately on: (i) the original dataset $\mathcal{D}_{1}$, (ii) the paraphrased dataset $\mathcal{D}_{2}$ generated using GPT-4o-mini, and (iii) the combined dataset $\mathcal{D}_{1}\cup\mathcal{D}_{2}$. Evaluation was performed across all possible test configurations; original, paraphrased, and combined to assess within-domain and cross-domain generalization.
The core objective is to test the causal invariance hypothesis, which posits that augmentations preserving causal semantics ($P(y|x^{\prime})=P(y|x)$) encourage models to learn representations invariant to superficial linguistic variations. Formally, we aim to minimize the expected difference in accuracy across domains:
$\underset{\theta}{\text{argmin}}\;\mathbb{E}_{d_{i},d_{j}\in\{\mathcal{D}_{1},\mathcal{D}_{2}\}}\big[\,|\text{Acc}_{d_{i}}(h_{\theta})-\text{Acc}_{d_{j}}(h_{\theta})|\,\big],$ (11)
where $h_{\theta}$ denotes the trained classifier. A smaller cross-domain difference indicates stronger invariance to domain shift induced by paraphrastic augmentation.
4.1.1 Results on Yoruba
Tables 2 and 3 summarize the validation and test accuracies for Yoruba, respectively. The combined configuration consistently yields higher accuracy across domains, confirming that integrating both original and paraphrased data enhances generalization.
Table 2: Validation results on Yoruba subset (Logistic Classifier).
Train Dataset Test Dataset Accuracy
Original Original 0.6787
Original Paraphrased 0.6602
Original Combined 0.6925
Paraphrased Original 0.6584
Paraphrased Paraphrased 0.6787
Paraphrased Combined 0.6907
Combined Original 0.6999
Combined Paraphrased 0.6898
Combined Combined 0.6934
Table 3: Test results on Yoruba subset (Logistic Classifier).
Train Dataset Test Dataset Accuracy
Original Test Set 0.6799
Paraphrased Test Set 0.6747
Combined Test Set 0.6892
Across the Yoruba results, the model trained on the combined dataset achieved the highest test accuracy of 68.92%, compared to 67.99% for the original-only model and 67.47% for the paraphrased-only model. This demonstrates that exposure to both natural and paraphrased data improves robustness to distributional variation. Moreover, the smaller cross-domain deviation in validation accuracies ($|\text{Acc}_{\text{orig}}-\text{Acc}_{\text{para}}|\approx 0.01$) further confirms increased invariance to paraphrastic shifts.
4.1.2 Results on Igbo
Tables 4 and 5 present the corresponding results for the Igbo subset. The general tend in this case, does not mirror Yoruba. While the combined dataset achieves stable performance across domains, improvements are more pronounced on our original data due to a larger semantic drift between original and paraphrased data.
Table 4: Validation results on Igbo subset (Logistic Classifier).
Train Dataset Test Dataset Accuracy
Original Original 0.7590
Original Paraphrased 0.6429
Original Combined 0.7230
Paraphrased Original 0.6922
Paraphrased Paraphrased 0.6834
Paraphrased Combined 0.7502
Combined Original 0.7370
Combined Paraphrased 0.6895
Combined Combined 0.7326
Table 5: Test results on Igbo subset (Logistic Classifier).
Train Dataset Test Dataset Accuracy
Original Test Set 0.7283
Paraphrased Test Set 0.7084
Combined Test Set 0.7176
While the Igbo combined model achieved a slightly lower test accuracy (71.76%) compared to the original-only model (72.83%), it exhibited more consistent validation performance across paraphrastic and combined domains, indicating reduced overfitting to surface lexical cues. This suggests that the degree of causal alignment between paraphrased and original text affects the strength of invariance gained.
Overall, results confirm that causal-preserving data augmentation yields measurable improvements in cross-domain robustness. In Yoruba, where paraphrases were semantically well-aligned, the combined model achieved the best generalization with the smallest invariance gap $\Delta_{\text{inv}}$. In Igbo, where GPT-4o paraphrases introduced a larger distributional shift, the combined model still maintained competitive performance, albeit with smaller gains.
These findings support the hypothesis that training on both original and paraphrased data enhances model invariance to superficial linguistic variability. The modest but consistent improvements, particularly in Yoruba, shows the potential of causal data augmentation as a lightweight strategy for generalization in low-resource sentiment classification.
4.2 Invariance via Causal Representation Learning
4.2.1 Training Configuration
We evaluate three transformer-based models: XLM-R, a multilingual RoBERTa model trained on 100 languages[30] ; Afro-XLMR Large, a specialized version fine-tuned for African languages[26] and Afro-XLMR Large-76L[31], an extended variant with additional African language coverage. For each model, we compare two training configurations: CT=False, which implements standard training where aspect and review are concatenated, and CT=True, which employs causal training where aspect and review are treated separately.
All models were fine-tuned with consistent hyperparameters including a learning rate of 2e-5, batch size of 16, maximum sequence length of 128, 10 training epochs, weight decay of 0.01, and the AdamW optimizer. For CT=True models, we implemented DINER’s causal inference components including backdoor adjustment with $\tau=0.1$, counterfactual reasoning using shared encoders, and TIE calculation with ensemble fusion function.
We evaluate performance using accuracy, macro F1-score, and Aspect Robustness Score (ARS), which measures performance under aspect perturbations[4]. To assess out-of-domain (OOD) generalization, we employ cross-language evaluation by training on one language and testing on others, as well as perturbation tests including RevTgt (reversing the polarity of target aspects), RevNon (reversing the polarity of non-target aspects), and AddDiff (adding aspects with different polarity).
4.2.2 Results on Convergence Analysis
A key finding is that causal training (CT=True) typically converges faster across multiple transformer models. As shown in our accuracy charts (Figures 3(a)-3(c)), models trained with causal inference require fewer steps to reach comparable performance. XLM-R with CT=True converges 13 steps earlier than CT=False, while Afro-XLMR Large with CT=True shows steeper initial learning curves, and Afro-XLMR Large-76L exhibits similar patterns with accelerated convergence. This efficiency advantage is particularly valuable in resource-constrained African AI contexts, enabling faster training cycles and reduced computational demands.
(a) XLM-R accuracy
(b) Afro-XLMR-large accuracy
(c) Afro-XLMR-large-76L accuracy
Figure 3: Comparison of accuracy charts. (a) XLM-R, (b) Afro-XLMR-large, (c) Afro-XLMR-large-76L.
4.2.3 Results on Performance Across Languages
Performance varies significantly across African languages, reflecting both linguistic differences and translation quality. Table 6 shows that for the 5 low-resource languages studied in depth, results differ considerably. Afrikaans demonstrates the highest overall performance, likely due to its similarity to English and high-quality translations. Amharic shows the lowest performance across all models, reflecting script and morphological complexity. Hausa exhibits moderate performance with significant gains under causal training. Igbo shows notable improvement with CT=True in XLM-R but less pronounced gains in AfroXLM variants. Yoruba follows a similar pattern to Igbo, with causal training providing modest improvements. These results highlight the language-dependent nature of causal inference benefits, with some languages benefiting more than others.
Table 6: Performance on test set for 5 low-resource languages (Accuracy %)
Model Setting Afrikaans Amharic Hausa Igbo Yoruba
XLM-R CT=False 78.2 62.1 70.4 71.5 72.3
XLM-R CT=True 79.6 63.5 72.8 74.2 75.1
Afro-XLMR Large CT=False 80.4 64.7 72.5 73.7 74.6
Afro-XLMR Large CT=True 81.1 65.2 74.1 74.9 75.8
Afro-XLMR Large-76L CT=False 81.6 65.3 73.2 74.5 75.4
Afro-XLMR Large-76L CT=True 82.5 66.1 74.9 76.1 77.2
4.2.4 OOD Generalization Results
The theoretical advantage of causal models in OOD settings shows mixed results in practice. Table 7 reveals that XLM-R shows better generalization with CT=True for Afrikaans, Igbo, and Yoruba, while AfroXLM-R Large underperforms with CT=True on many OOD tasks, though exceptions exist where causal training significantly improves OOD performance. These inconsistencies likely stem from translation noise affecting causal variable alignment, model architecture sensitivity to causal training, and varying degrees of distributional shift across languages.
Table 7: Performance in OOD settings (Accuracy %)
Model Setting Afrikaans Amharic Hausa Igbo Yoruba
XLM-R CT=False 52.4 48.3 49.2 47.6 48.5
XLM-R CT=True 55.8 47.9 48.7 51.2 52.1
Afro-XLMR Large CT=False 54.7 50.1 51.3 49.8 49.6
Afro-XLMR Large CT=True 53.5 48.7 49.5 48.4 50.3
Afro-XLMR Large-76L CT=False 41.02 41.14 41.28 37.79 39.41
Afro-XLMR Large-76L CT=True 34.52 34.92 35.13 30.89 33.41
4.2.5 Case Studies
Table 8 presents illustrative examples demonstrating the robustness improvements from causal training. In the original example, both baseline and DINER correctly classify positive sentiment. For RevTgt cases, both models correctly adjust when target sentiment is reversed. However, for RevNon scenarios, DINER correctly maintains focus on target aspect despite negative non-target aspects, while baseline models fail. Similarly, with AddDiff examples, DINER shows robustness when additional negative aspects are introduced. These examples highlight DINER’s ability to isolate target aspect sentiment from confounding context—a crucial capability for practical ABSA applications.
Table 8: Example case studies. The corresponding gold labels and the predictions for each example are presented.
Type Examples (Target Aspect: food) Gold Baseline DINER
Original The food is top notch, the service is attentive, and the atmosphere is great. Positive Positive ✓ Positive ✓
RevTgt The food is nasty, but the service is attentive, and the atmosphere is great. Negative Negative ✓ Negative ✓
RevNon The food is top notch, the service is heedless, but the atmosphere is not great. Positive Negative ✗ Positive ✓
AddDiff The food is top notch, the service is attentive, and the atmosphere is great, but music is too heavy, waiters is angry and staff is arrogant. Positive Negative ✗ Positive ✓
5 Conclusion and Future Work
In this study, we systematically investigated the efficacy of two distinct causal domain generalization techniques in the challenging context of low-resource African languages. By examining both a data-level intervention through Causal Data Augmentation (CDA) and a representation-level intervention with Invariant Causal Representation Learning (ICRL), we aimed to understand their complementary strengths and practical trade-offs.
Our findings confirm that both approaches offer tangible benefits for enhancing model robustness against distribution shifts. The CDA experiments on the NaijaSenti corpus demonstrated that augmenting training data with semantically equivalent paraphrases is a lightweight yet effective strategy for improving cross-domain generalization. The positive results in Yoruba, where the combined model achieved higher accuracy and a smaller invariance gap, underscore the potential of this method when paraphrase quality is high. Conversely, the results in Igbo highlight a key dependency: the success of CDA is contingent on the semantic fidelity of the generated counterfactuals, as significant distributional drift can temper performance gains.
The application of the DINER framework to our newly created Afri-SemEval benchmark provided a more nuanced picture. While causal training consistently accelerated model convergence and improved in-domain accuracy for many of the 17 languages, its impact on Out-of-Distribution (OOD) generalization was varied. This inconsistency reveals a critical insight: the theoretical promise of sophisticated causal models can be attenuated by the practical realities of cross-lingual transfer. We identify a "dual-shift" problem, where models must contend not only with the intended linguistic domain shift but also with an unintended, artifact-based shift introduced by translation noise and aspect alignment errors. This suggests that the language-dependent nature of the observed gains is a direct reflection of this complex interaction.
Looking ahead, our work points to several promising research directions. First, the limitations of our heuristic-based aspect alignment highlight the need for more robust methods to adapt complex models like DINER to typologically diverse languages. Future work could explore jointly trained alignment models or the development of syntactic parsers tailored to African languages. Second, the variable quality of LLM-generated paraphrases suggests a need for more controlled generation techniques or human-in-the-loop validation pipelines to maximize the benefits of CDA. Finally, the "dual-shift" phenomenon calls for the development of novel domain generalization methods that are explicitly designed to be robust not only to natural linguistic variation but also to the systematic noise inherent in cross-lingual transfer. A hybrid approach, perhaps using CDA to create more robust pre-trained models before applying representation-level debiasing, could prove to be a fruitful avenue for future exploration.
Acknowledgments:
We are grateful to ML Collective for generously providing the computational resources required for this study. The success of our extensive experiments, particularly the resource-intensive large language model (LLM) fine-tuning and the multi-language DINER training, was directly enabled by their dedicated compute support.