### v2

# When Distributions Shifts: Causal Generalization for Low-Resource Languages

Abstract
Machine learning models often fail under distribution shifts, a problem exacerbated in low-resource settings where limited data restricts robust generalization. Domain generalization (DG) methods address this challenge by learning representations that remain invariant across domains, frequently leveraging causal principles. In this work, we study two causal DG approaches for low-resource natural language processing. First, we apply causal data augmentation using GPT-4o-mini to generate counterfactual paraphrases for sentiment classification on the NaijaSenti Twitter corpus in Yoruba and Igbo. Second, we investigate invariant causal representation learning with the Debiasing in Aspect Review (DINER) framework for aspect-based sentiment analysis. We extend DINER to a multilingual setting by introducing Afri-SemEval, a dataset of 17 languages translated from SemEval-2014 Task. Experiments show improved robustness to unseen domains, with consistent gains from counterfactual augmentation and enhanced out-of-distribution performance from causal representation learning across multiple languages.
Keywords Causal representation $\cdot$ Low-resource $\cdot$ Data augmentation $\cdot$ Language translation $\cdot$ LLMs
1 Introduction
Machine learning models have achieved notable success in natural language processing, but their effectiveness relies on the assumption that training and test data are identically distributed. In practice, data distributions shift across time, geography, and linguistic contexts, leading to performance degradation on unseen domains. This challenge is severe in low-resource settings. In a study conducted by Adebara et al. [1], they introduced the “Sahara” benchmark for African NLP and showed that out of 517 African languages in their benchmark, only 45 languages were present in more than one dataset, with most appearing only in basic language identification tasks. Such extreme data sparsity and fragmentation make it infeasible to rely on scale alone, amplifying the impact of distribution shifts and the need for domain generalization methods that can learn invariant representations under limited and heterogeneous data conditions [2, 3, 4, 5].
In machine learning, domain generalization (DG) addresses the challenge of performing well on unseen domains by learning representations that remain invariant across multiple source domains without access to the target distribution. This DG formulation and its associated solutions also extend to low-resource settings, where data scarcity makes robustness to distribution shifts especially important. Recent DG approaches increasingly emphasize learning stable and invariant features across domains; causal perspectives have been shown to be one effective way to formalize and motivate such invariance under domain shifts [6].
In this study, we examine two complementary causal DG approaches for low-resource sentiment analysis: Causal Data Augmentation (CDA), which generates sentiment-preserving paraphrases to simulate controlled domain shifts, and Causal Representation Learning (CRL), which applies structural causal models with backdoor adjustment and counterfactual reasoning to isolate causal sentiment paths. Figure 1 illustrates our unified framework. For CDA, we use the Yoruba and Igbo subsets of NaijaSenti [7]. For CRL, we introduce Afri-SemEval, a new benchmark translating SemEval-2014 Task 4 [8] into 17 African languages.
Figure 1: Overview of our unified causal framework. Left (CDA): Paraphrased data is generated via GPT-4o-mini under the constraint $P(Y|X^{\prime})=P(Y|X)$, then combined with original data for invariant representation learning. Right (CRL): The DINER mechanism applies aspect-context alignment, backdoor adjustment, and counterfactual reasoning to produce debiased sentiment predictions.
Our contributions are: (1) a unified framework combining data-level and representation-level causal interventions for low-resource sentiment analysis; (2) Afri-SemEval, a multilingual ABSA benchmark covering 17 African languages; and (3) empirical evidence that causal mechanisms enhance generalization under distribution shifts in low-resource multilingual NLP.
In the rest of this paper, Section 2 reviews related work on domain generalization and causal NLP. Section 3 details our methodology for both CDA and CRL approaches. Section 4 presents experimental results and analysis, and Section 5 concludes the paper, discusses limitations and future directions.
2 Related Work
Domain generalization (DG) in NLP aims to improve model robustness under distribution shifts.
Table 1 summarizes key Domain Generalization approaches that have been explored by researchers, highlighting whether they use Causal Data Augmentation (CDA) or Causal Representation Learning (CRL) and their main contributions and limitations.
While prior work [9] have explored causal data augmentation (CDA) for low-resource languages, controlled sentiment-preserving paraphrasing for simulating causal domain shifts is largely unexplored. Additionally, causal representation learning (CRL) has mostly been applied to high-resource languages.
Our study explores two causal frameworks for low-resource African languages; CDA and CRL. Unlike previous works, we explored CDA via sentiment-preserving paraphrases in Yoruba and Igbo. For CRL, we introduced Afri-SemEval, a multilingual ABSA dataset covering 17 African languages, enabling evaluation of domain generalization under distribution shifts in low-resource multilingual settings.
Authors Technique Research Goal CDA/CRL Key Contribution
Wu et al. [10] Multilingual ABSA Benchmark (M-ABSA) Multilingual aspect-based sentiment analysis CRL Addressed data scarcity via benchmark construction and fine-tuning of generative models.
Muhammad et al. [7]; Muhammad et al. [11] Dataset Creation (NaijaSenti, AfriSenti) Sentiment analysis for African languages CDA Pioneered African language sentiment datasets but limited to thousands of examples.
Ibrahim et al. [12]; Maqsood et al. [13] Low-resource Benchmark Creation Sentiment analysis for Hausa and Urdu CRL Introduced language-specific benchmarks that improve coverage but not address robustness in domain shifts.
Kaushik et al. [14] Counterfactual Augmentation Reduce spurious feature sensitivity CDA Showed 40–60% reduction in spurious correlations via counterfactuals.
Šmíd et al. [15] Constrained Decoding Syntactic validity in cross-lingual generation CRL Enforced syntactic constraints during generation without addressing semantic or causal robustness.
Balashankar et al. [16] Active Learning + Counterfactual Generation Efficient robustness with minimal annotation CDA Achieved 18–20% robustness gains with 10% human annotation.
Møller et al. [9] GPT-4 Data Augmentation Low-resource classification tasks CDA Noted quality degradation for low-resource languages.
Wu et al. [17] DINER Debias ABSA via causal inference CRL Backdoor adjustment and counterfactual reasoning to English ABSA.
Table 1: Overview of domain generalization approaches relevant to low-resource NLP. Techniques are categorized by authors, method, research goal, and type of causal intervention: Causal Data Augmentation (CDA) or Causal Representation Learning (CRL). Key contributions highlight each method’s main achievements and limitations.
3 Methodology
This section presents the methodological design adopted in this study, which comprises of two experimental frameworks: (1)Invariance via Causal Data Augmentation and (2) Invariance via Causal Representation Learning.
3.1 Invariance via Causal Data Augmentation
Here, we investigate causal data augmentation as a strategy for improving sentiment classification in low-resource African languages through representation invariance. The approach is grounded in the principles of causal representation learning: if the sentiment label $y$ depends on the semantic content of a tweet $x$, then augmentations that perturb only non-causal linguistic features (e.g., surface form, style, or syntax) should preserve the conditional probability distribution $P(y|x)$.
We used the Yoruba and Igbo subsets of the NaijaSenti corpus[7], a multilingual sentiment twitter dataset covering four Nigerian languages: Yoruba, Igbo, Hausa, and Nigerian Pidgin. Each instance is a tweet–label pair $(x_{i},y_{i})$, where $x_{i}$ represents the tweet and $y_{i}\in\{\text{positive},\text{negative},\text{neutral}\}$ is its sentiment label.
3.1.1 Experiment Procedure
Notation.
Let $\mathcal{X}$ denote the input text space and $\mathcal{Y}$ the sentiment label space. Let $\mathcal{D}_{1}=\{(x_{i},y_{i})\}_{i=1}^{N}$ be the original sentiment dataset sampled from a joint distribution $P(X,Y)$, where $x_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{Y}$. Let $f:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ be a paraphrasing function producing $x^{\prime}=f(x)$. We denote the paraphrased dataset as $\mathcal{D}_{2}=\{(x^{\prime}_{i},y_{i})\}_{i=1}^{N}$, and the combination of the paraphrased and original dataset as $\mathcal{D}_{\text{comb}}=\mathcal{D}_{1}\cup\mathcal{D}_{2}$.
Data Augmentation via Paraphrasing.
To simulate controlled domain shifts while preserving label semantics, we construct our paraphrased dataset $\mathcal{D}_{2}$ using a sentiment-preserving augmentation function $f:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ such that $P(Y\mid X^{\prime})=P(Y\mid X)$. The paraphrases are generated using the GPT-4o-mini model [18] with a carefully designed prompt that explicitly enforces sentiment and semantic preservation while introducing lexical and syntactic variation. The prompt template is shown below:
“You have been provided with a {language} tweet. Paraphrase ‘{tweet}’ in a way that preserves its sentiment and meaning but changes its wording. Return only the paraphrased tweet.”
This process yields a paraphrased dataset $\mathcal{D}_{2}$ that is semantically aligned with the original dataset $\mathcal{D}_{1}$. We observed that GPT-4o-mini preserves sentiment polarity while introducing lexical and syntactic diversity which retains the underlying sentiment and lies in the same semantic neighborhood. The illustration in Table 2 are examples of original tweets and their corresponding paraphrases (‘Y” denotes Yoruba, ‘I” denotes Igbo).
Original Tweet ($X$) Paraphrased Tweet ($X^{\prime}$) (GPT-4o-mini) Label ($Y$)
Y: “Oúnjẹ yìí dára púpọ̀, mo fẹ́ràn rẹ̀ gan-an” “Oúnjẹ náà dùn púpọ̀, ó wù mí gidigidi” positive
I: “Ahịa a jọrọ njọ, ọ gaghị m azụta ya ọzọ” “Ngwa ahịa a adịghị mma, agaghị m alọghachikwuru ya” negative
Y: “Ayé ilá ilá ikàn wọ ẹ̀wù ẹ̀jẹ̀ sọ ọrún ọ̀pá gbédìí fáyé” “Ayé ilá ilá ikàn wọ aṣọ ẹ̀jẹ̀ sọ ọrún ọ̀pá gbédìí fáyé” negative
Table 2: Examples of sentiment-preserving paraphrases generated by GPT-4o-mini for Yoruba and Igbo tweets.
One of the main challenges in guiding LLMs to generate diverse variants while maintaining key information arises for proverbial and sarcastic tweets, where sentiment is often encoded through more nuanced expressions. In such cases, the model’s ability to retain variations of the original sentiment without diluting the underlying proverbial meaning or sarcasm is limited. To mitigate this issue, we employ a human annotation and correction mechanism for a subset of tweets exhibiting these challenges.
Our resulting output exhibits a shift in the marginal input distribution:
$\begin{split}P_{\mathcal{D}_{2}}(X^{\prime})&\neq P_{\mathcal{D}_{1}}(X),\\ P_{\mathcal{D}_{2}}(Y\mid X^{\prime})&\approx P_{\mathcal{D}_{1}}(Y\mid X).\end{split}$ (1)
Figure 2: Flow of $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{\text{comb}}$
For each language subset, we train three logistic regression classifiers: Original-only, trained on $\mathcal{D}_{1}$; Paraphrased-only, trained on $\mathcal{D}_{2}$; and Combined, trained on $\mathcal{D}_{\text{comb}}$ with a balanced 50–50 mix (this 50:50 mixing strategy is used to avoid bias stemming from over-representation of either original or paraphrased samples during training). We choose logistic regression for its interpretability and well-understood optimization properties, allowing us to isolate the effects of data augmentation and distributional shift without confounding factors introduced by model complexity. To assess robustness under the introduced distribution shift, we adopt a cross-domain evaluation protocol. Each model is evaluated on $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{\text{comb}}$, enabling systematic comparison of in-domain and cross-domain generalization performance. See Figure 3.
Each classifier $h_{\theta}$ is trained by minimizing the empirical cross-entropy loss:
$\hat{\mathcal{L}}(\theta)=\frac{1}{|\mathcal{D}_{\text{train}}|}\sum_{(x,y)\in\mathcal{D}_{\text{train}}}\ell\big(h_{\theta}(g_{\psi}(\phi(x))),y\big),$ (2)
where $\phi$ denotes text preprocessing (tokenization, normalization, stopword removal), and $g_{\psi}$ is a fixed sentence embedding function implemented using the paraphrase-multilingual-MiniLM-L12-v2 model [19]. Our model performance was measured using “accuracy” for classification tasks. This experimental design enables a principled empirical test of our causal invariance hypothesis, by evaluating whether models trained with label-preserving augmentations improves generalization under distributional shift in low-resource settings.
Figure 3: Flow of the experiment procedure used in Invariance via Causal Data Augmentation.
3.2 Invariance via Causal Representation Learning
Invariance via Causal Representation Learning (CRL) during inference is one of the causal domain generalization techniques we explore in low-resource settings. To achieve this, we adopt the DINER: Debiasing Aspect-based Sentiment Analysis with Multi-variable Causal Inference [17], a debiasing approach for the Aspect-Based Sentiment Analysis (ABSA) task. We discuss our methodology in what follows.
3.2.1 Debiasing in Aspect Review (DINER)
Recent studies have proposed several debiasing strategies, including argumentation-based methods [20], reweighting-based training [21], and causal inference-based approaches [22, 23]. Among these, causal inference has attracted significant interest for its solid theoretical foundation and minimal disruption to existing learning paradigms. See Section A for more details on debiasing in aspect review using DINER.
3.2.2 Data
We introduce Afri-SemEval, a multilingual dataset created by translating the SemEval 2014 dataset into 17 African languages to support research on domain generalization in low-resource contexts. Figure 4 illustrates our dataset creation pipeline. We leverage the Microsoft Azure Translator API for one-to-many translation, where English reviews are batch-processed with configurations settings to comply with API constraints and translated simultaneously into all target languages. The pipeline parses JSON responses, maps translations to source review IDs, and applies a semantic similarity heuristic to align aspect terms across languages. The dataset spans diverse languages from multiple African language families, as shown in Table 3.
Figure 4: Afri-SemEval dataset creation pipeline showing the translation workflow for SemEval-2014 to 17 African languages using Azure API.
S/N Language Code S/N Language Code
1 Afrikaans af 10 Kinyarwanda rw
2 Amharic am 11 Shona sn
3 Hausa ha 12 Somali so
4 Igbo ig 13 S. Sotho st
5 Lingala ln 14 Tswana tn
6 Luganda lug 15 Xhosa xh
7 N. Sotho nso 16 Yoruba yo
8 Chichewa nya 17 Zulu zu
9 Kirundi run
Table 3: List of 17 African supported languages
Each translated instance maintains the original SemEval-2014 format, comprising review text, target aspect, and sentiment polarity (positive, negative, or neutral). The translation process involved ensuring aspect alignment, polarity preservation, and accounting for linguistic and cultural nuances in sentiment expression.
A major challenge in adapting DINER to African languages lies in aspect term alignment. The original framework assumes a direct mapping between aspect terms and their positions in the review, but translation often disrupts this due to word order variation, morphological complexity (e.g., noun classes in Bantu languages), and differences in expression length (multiword to single-word mappings, and vice versa).
We address this with an alignment heuristic based on semantic similarity. Embeddings are computed for translated aspect terms and all n-grams in the review. The n-gram with the highest similarity to the aspect term is then selected as its aligned position for DINER processing. While effective for single-token aspects, this heuristic remains challenging for multi-token and non-contiguous aspect terms.
4 Experiment and Result
4.1 Invariance via Causal Data Augmentation
All experiments were conducted on the Yoruba and Igbo subsets of the NaijaSenti dataset[7], following the causal data augmentation setup described in Section 3.1. For each language, three logistic regression classifiers were trained separately on: (i) the original dataset $\mathcal{D}_{1}$, (ii) the paraphrased dataset $\mathcal{D}_{2}$ generated using GPT-4o-mini, and (iii) the combined dataset $\mathcal{D}_{1}\cup\mathcal{D}_{2}$. Evaluation was performed across all possible test configurations; original, paraphrased, and combined to assess within-domain and cross-domain generalization.
The core objective is to test the causal invariance hypothesis, which posits that augmentations preserving causal semantics ($P(y|x^{\prime})=P(y|x)$) encourage models to learn representations invariant to superficial linguistic variations. Formally, we aim to minimize the expected difference in accuracy across domains:
$\underset{\theta}{\text{argmin}}\;\mathbb{E}_{d_{i},d_{j}\in\{\mathcal{D}_{1},\mathcal{D}_{2}\}}\big[\,|\text{Acc}_{d_{i}}(h_{\theta})-\text{Acc}_{d_{j}}(h_{\theta})|\,\big],$ (3)
where $h_{\theta}$ denotes the trained classifier.
4.1.1 Results on Yoruba
Tables 4 and 5 summarize the validation and test accuracies for Yoruba, respectively. The combined configuration consistently yields higher accuracy across domains, confirming that integrating both original and paraphrased data enhances generalization.
Train Dataset Test Dataset Accuracy
Original Original 0.6787
Original Paraphrased 0.6602
Original Combined 0.6925
Paraphrased Original 0.6584
Paraphrased Paraphrased 0.6787
Paraphrased Combined 0.6907
Combined Original 0.6999
Combined Paraphrased 0.6898
Combined Combined 0.6934
Table 4: Validation results on Yoruba subset.
Train Dataset Test Dataset Accuracy
Original Test Set 0.6799
Paraphrased Test Set 0.6747
Combined Test Set 0.6892
Table 5: Test results on Yoruba subset.
Across the Yoruba results, the model trained on the combined dataset achieved the highest test accuracy of 68.92%, compared to 67.99% for the original-only model and 67.47% for the paraphrased-only model. This demonstrates that exposure to both natural and paraphrased data improves robustness to distributional variation. Moreover, the smaller cross-domain deviation in validation accuracies ($|\text{Acc}_{\text{orig}}-\text{Acc}_{\text{para}}|\approx 0.01$) further confirms increased invariance to paraphrastic shifts.
4.1.2 Results on Igbo
Tables 6 and 7 present the corresponding results for the Igbo subset. The general trend in this case, does not mirror Yoruba. While the combined dataset achieves stable performance across domains, improvements are more pronounced on our original data due to a larger semantic drift between original and paraphrased data.
Train Dataset Test Dataset Accuracy
Original Original 0.7590
Original Paraphrased 0.6429
Original Combined 0.7230
Paraphrased Original 0.6922
Paraphrased Paraphrased 0.6834
Paraphrased Combined 0.7502
Combined Original 0.7370
Combined Paraphrased 0.6895
Combined Combined 0.7326
Table 6: Validation results on Igbo subset.
Train Dataset Test Dataset Accuracy
Original Test Set 0.7283
Paraphrased Test Set 0.7084
Combined Test Set 0.7176
Table 7: Test results on Igbo subset.
While the Igbo combined model achieved a slightly lower test accuracy (71.76%) compared to the original-only model (72.83%), it exhibited more consistent validation performance across paraphrastic and combined domains, indicating reduced overfitting to surface lexical cues. This suggests that the degree of causal alignment between paraphrased and original text affects the strength of invariance gained.
Overall, our results confirm that causal-preserving data augmentation yields measurable improvements in cross-domain robustness. In Yoruba, where paraphrases were semantically well-aligned, the combined model achieved the best generalization with the smallest invariance gap $\Delta_{\text{inv}}$. In Igbo, where paraphrases had a larger distributional shift, the combined model still maintained competitive performance, with smaller gains. These findings suggests that training on both original and paraphrased data enhances model invariance to superficial linguistic variability. The modest but consistent improvements show the potential of causal data augmentation as a lightweight strategy for generalization in low-resource sentiment classification.
4.2 Invariance via Causal Representation Learning
For our experiments, we evaluated three transformer-based models: XLM-R [24], Afro-XLMR Large [25], and Afro-XLMR Large-76L [26]. For each model, we compare two training configurations: CT=False (standard training with concatenated aspect and review) and CT=True (causal training with separate aspect and review processing).
All models were fine-tuned with learning rate 2e-5, batch size 16, maximum sequence length 128, 10 epochs, weight decay 0.01, and AdamW optimizer. For CT=True models, we implemented DINER’s causal inference components including backdoor adjustment ($\tau=0.1$), counterfactual reasoning, and TIE calculation with ensemble fusion.
We use accuracy as metrics to access the performance of our approach. To assess out-of-distribution (OOD) generalization, we employ cross-language evaluation and perturbation tests including RevTgt (reversing target aspect polarity), RevNon (reversing non-target aspect polarity), and AddDiff (adding aspects with different polarity).
4.2.1 Results
A key finding is that causal training (CT=True) typically converges faster across multiple transformer models. As shown in Figures 5, 6, and 7, models trained with causal inference require fewer steps to reach comparable performance. This efficiency advantage is particularly valuable in resource-constrained contexts, enabling faster training cycles and reduced computational demands.
Figure 5: XLM-R accuracy comparison showing convergence patterns for CT=False and CT=True configurations.
Figure 6: Afro-XLMR-large accuracy comparison showing convergence patterns for CT=False and CT=True configurations.
Figure 7: Afro-XLMR-large-76L accuracy comparison showing convergence patterns for CT=False and CT=True configurations.
Performance varies across African languages, reflecting both linguistic differences and translation quality. Table 8 presents the performance results across five low-resource African languages for three transformer-based models under both standard training (CT=False) and causal training (CT=True) configurations. Afrikaans demonstrates the highest overall performance, likely due to its similarity to English. Amharic shows the lowest performance, reflecting script and morphological complexity. Hausa, Igbo, and Yoruba exhibit moderate performance with varying gains under causal training.
Model Setting Afrikaans Amharic Hausa Igbo Yoruba
XLM-R CT=False 78.2 62.1 70.4 71.5 72.3
XLM-R CT=True 79.6 63.5 72.8 74.2 75.1
Afro-XLMR Large CT=False 80.4 64.7 72.5 73.7 74.6
Afro-XLMR Large CT=True 81.1 65.2 74.1 74.9 75.8
Afro-XLMR Large-76L CT=False 81.6 65.3 73.2 74.5 75.4
Afro-XLMR Large-76L CT=True 82.5 66.1 74.9 76.1 77.2
Table 8: Performance on test set for 5 low-resource languages (Accuracy %)
The theoretical advantage of causal models in OOD settings shows mixed results in practice. Table 9 shows the OOD generalization results across the same five languages, demonstrating varying effectiveness of causal training across different model architectures. The results reveals that XLM-R generally benefits from CT=True for Afrikaans, Igbo, and Yoruba, while AfroXLM-R Large shows inconsistent patterns. These inconsistencies likely stem from translation noise affecting causal variable alignment and varying degrees of distributional shift across languages.
Model Setting Afrikaans Amharic Hausa Igbo Yoruba
XLM-R CT=False 52.4 48.3 49.2 47.6 48.5
XLM-R CT=True 55.8 47.9 48.7 51.2 52.1
Afro-XLMR Large CT=False 54.7 50.1 51.3 49.8 49.6
Afro-XLMR Large CT=True 53.5 48.7 49.5 48.4 50.3
Afro-XLMR Large-76L CT=False 41.02 41.14 41.28 37.79 39.41
Afro-XLMR Large-76L CT=True 34.52 34.92 35.13 30.89 33.41
Table 9: Performance in OOD settings (Accuracy %)
In section  A.3, we further discussed some examples and case studies that demonstrates the robustness improvements from the causal training in our CRL experiments. See table  10.
5 Conclusion and Future Work
In this study, we investigated two causal domain generalization techniques for low-resource African languages: Causal Data Augmentation (CDA) and Invariant Causal Representation Learning (ICRL) via the DINER framework.
Our findings confirm that both approaches enhance model robustness against distribution shifts. CDA experiments on NaijaSenti showed that augmenting with semantically equivalent paraphrases improves cross-domain generalization, particularly in the Yoruba dataset where paraphrase quality was high. In contrast, Igbo results highlight that success depends on semantic fidelity, as distributional drift can limit gains. Applying DINER to the Afri-SemEval dataset demonstrated that causal training accelerates convergence and improves in-domain accuracy across 17 languages, while out-of-distribution generalization varied. This variability reveals a dual-shift problem, where models face both linguistic domain shifts and artifacts introduced by translation noise and aspect misalignment.
Future work should address three directions: developing robust aspect alignment methods for typologically diverse languages, improving LLM-generated paraphrase quality through controlled generation or human-in-the-loop validation, and designing domain generalization methods robust to both natural linguistic variation and cross-lingual transfer noise. A hybrid approach combining CDA-enhanced pretraining with representation-level debiasing may be effective.
Ethical Considerations This work focuses on low-resource languages, which are historically underrepresented in NLP. While our dataset and models aim to improve language coverage and model generalization in distribution shift settings, they may still reflect cultural or social biases present in the source data.
We only use publicly available text and synthetic paraphrases, avoiding private or sensitive information. Finally, generated paraphrases should respect cultural norms to avoid offensive or misleading outputs, ensuring responsible and ethical use of the resources we provide.
Limitations While this study demonstrates the effectiveness of causal domain generalization techniques in low-resource NLP, several limitations remain. First, the Afri-SemEval dataset relies on machine translation, which may introduce artifacts and fail to capture nuances of native African language expressions. Second, our experiments focus on sentiment analysis, and the generalizability of these findings to other NLP tasks remains unexamined. Third, the computational cost of counterfactual data augmentation using large language models may limit scalability in resource-constrained settings. Finally, our evaluation covers only 17 African languages, and broader linguistic coverage would strengthen the conclusions.
Acknowledgment:
We are grateful to ML Collective for generously providing the computational resources required for this study. The success of our extensive experiments, in particular, the resource-intensive large language model (LLM) fine-tuning and the multi-language DINER training, was directly enabled by their dedicated compute support.
Appendix A Appendix
A.1 DINER FRAMEWORK
The DINER framework models ABSA as a Structural Causal Model (SCM) with five key variables: confounder $C$ (prior contextual knowledge), review $R$, aspect $A$, fused knowledge $K$ (derived from $R$ and $A$), and sentiment label $L$. The causal relations are defined as $R\rightarrow K\leftarrow A$, $K\rightarrow L$, $R\rightarrow L\leftarrow A$, and $C\rightarrow R,C\rightarrow L$. The causal structure is illustrated in Figure 9, showing the directed acyclic graph (DAG) that represents the relationships between these variables. The objective of DINER is to isolate the causal path $K\rightarrow L$ while eliminating spurious correlations through $R\rightarrow L\leftarrow A$ and minimizing confounding through $C\rightarrow R$ and $C\rightarrow L$. For details on causal graph construction, see [17].
A.2 Causal Variables Discovery
Figure 8 demonstrates how high-capacity neural networks can easily learn spurious correlations arising from annotation bias. These biases cause models to capture misleading dependencies rather than true causal relationships. Due to their high representational power, neural models are particularly vulnerable to such biases, often learning shortcuts instead of underlying latent features ([27, 28, 29]). For instance, reversing the polarity of target aspects can result in up to a 20% drop in accuracy, highlighting the model’s reliance on these correlations. The figure also supports the observation that models frequently depend on spurious correlations between target and non-target aspects in sentiment datasets such as [8]. In the “Original” setting, both review-only and aspect-only models perform well; however, when these correlations are disrupted in the “RevTgt,” “RevNon,” and “AddDiff” settings, performance decreases or becomes balanced. This confirms that neural models often exploit dataset biases rather than learning genuine target-specific sentiment representations.
Figure 8: Demonstration of spurious correlations learned from annotation bias ([17]).
Figure 9: Causal structure of the DINER framework ([17]).
Type Examples (Target Aspect: food) Gold Baseline DINER
Original The food is top notch, the service is attentive, and the atmosphere is great. Positive Positive ✓ Positive ✓
RevTgt The food is nasty, but the service is attentive, and the atmosphere is great. Negative Negative ✓ Negative ✓
RevNon The food is top notch, the service is heedless, but the atmosphere is not great. Positive Negative ✗ Positive ✓
AddDiff The food is top notch, the service is attentive, and the atmosphere is great, but the music is too heavy and the staff is arrogant. Positive Negative ✗ Positive ✓
Table 10: Illustrative case studies showing model robustness under target and non-target sentiment perturbations.
To mitigate the indirect confounding between sentiment words in the review and the context, DINER employs a backdoor adjustment intervention that addresses the $C\rightarrow R\rightarrow L$ pathway. The intervention uses the do-calculus formulation:
$\begin{split}P(L|do(R))&=\sum_{c}P(L|R,C)P(C)\\ &=\sum_{c}\frac{P(L,R|C)P(C)}{P(R|C)}\end{split}$ (4)
The debiased review representation is computed as:
[…]
To address the direct correlation between aspect and label ($A\rightarrow L$), DINER employs counterfactual reasoning using the Total Indirect Effect (TIE):
$\begin{split}TIE_{a,r}&=TE_{a,r}-NDE_{r}-NDE_{a}+IE_{a,r}\\ &=TE_{a,r}-(NDE_{r}+NDE_{a})\\ &\quad(\text{since }IE_{a,r}=0)\end{split}$ (6)
Following [22], we define:
$\begin{split}TE_{a,r}&=L_{a,r,k}-L_{a^{*},r^{*},k^{*}}\\ NDE_{a}&=L_{a,r^{*},k^{*}}-L_{a^{*},r^{*},k^{*}}\\ NDE_{r}&=L_{a^{*},r,k^{*}}-L_{a^{*},r^{*},k^{*}}\end{split}$ (7)
Thus:
$\begin{split}TIE_{a,r}=L_{a,r,k}-L_{a^{*},r,k^{*}}\\ -L_{a,r^{*},k^{*}}+L_{a^{*},r^{*},k^{*}}\end{split}$ (8)
The final prediction is implemented through a model ensemble ([17]):
$\begin{split}L_{a,r,k}&=L(A\!=\!a,R\!=\!r^{\prime},K\!=\!k)\\ &=\Psi(\zeta_{a},\zeta_{r^{\prime}},\zeta_{k})\\ &=\zeta_{k}\!+\!\tanh(\zeta_{a})\!+\!\tanh(\zeta_{r^{\prime}})\end{split}$ (9)
Where $\zeta_{r^{\prime}}$ is the output of the debiased review-only branch, $\zeta_{a}$ is the output of the aspect-only branch, and $\zeta_{k}$ is the output of the fused knowledge branch.
A.3 Case Studies
Table 10 presents illustrative examples demonstrating the robustness improvements from our causal training. In the original example, both baseline and DINER correctly classify positive sentiment. For RevTgt cases, both models correctly adjust when target sentiment is reversed. However, for RevNon scenarios, DINER correctly maintains focus on target aspect despite negative non-target aspects, while baseline models fail. Similarly, with AddDiff examples, DINER shows robustness when additional negative aspects are introduced. These examples highlight DINER’s ability to isolate target aspect sentiment from confounding context—a crucial capability for practical ABSA applications.